Case Eastern Junior College awarded 26 varsity letters in crew ,15 in swimming and 16 in soccer . If awards went to 46 students and onlt 2 lettered in all sports how many students lettered in 2 of the 3 sports.
Answer by edjones(8007)

There are a total of 26+15+16=57 letters.
If you remove the 6 letters that 2 students got, you have 24 crew, 13 swim, 14 soccer and 44 students remaining.
24+13+14=51
51-44=7 won 2 letters
.
Ed